### v1

# Linearly-scalable and entropy-optimal learning of nonstationary and nonlinear manifolds

Abstract
Unsupervised extraction of relevant low-dimensional manifolds from high-dimensional data is in core of many data analysis problems. Common linear methods like the Principal Component Analysis (PCA) and related linear approaches scale linearly with the data statistics size $T$ - but frequently fail to extract the nonlinear or changing in time (nonstationary) low-dimensional manifolds. Common nonlinear methods, like the t-distributed Stochastic Neighbour Embedding (t-SNE), Uniform Manifold Approximation and Projection (UMAP) - as well as their kernelized and parametric extensions (e.g., neural networks for parametric UMAP) - scale as $\mathcal{O}(T^{2})$, or $\mathcal{O}(T\textrm{log}(T))$ in the best case, frequently fail extracting nonstationary manifolds, and can distort results by generating method-induced artefacts (e.g., reveal clusters that are actually not present in the data, or do not reveal clusters when they are present, some examples are provided). PCA-based clustering methods scale as $\mathcal{O}(T)$ in every iteration, and can be used to extract linear nonstationary manifolds, but fail in the situations when the manifolds are nonlinear (some examples are also provided). Here we propose an Entropy-Optimal Manifold Clustering (EOMC) as a metricised and entropy-regularized extension of PCA clustering - and show that it mitigates the problems of the existing tools even in very nonstationary and nonlinear situations, while pertaining the favourable $\mathcal{O}(T)$ iteration complexity scaling. In comparison with the state-of-the-art linear and nonlinear methods on a set of noisy high-dimensional synthetic benchmarks, EOMC is demonstrated to provide a very robust artefact-free learning and reconstruction of low-dimensional manifolds from noisy, nonlinear and nonstationary data. Application to the Lorenz-96 dynamical system - a very popular model of a turbulent behaviour in one dimension - in chaotic and strongly-chaotic regimes reveals that its dynamics is essentially described by a metastable regime-switching process, making infrequent transitions between the very persistent three-dimensional attractive manifolds. The dimensionality of these manifolds appears to remain unchanged, and their overall number gradually grows with the growing external forcing of the Lorenz-96 model. At the same time, the Markovian mean exit times and relaxation times (that bound the predictability horizons for the identified regime-switching process) appear to decrease only very slowly with the growing external forcing - indicating approximately two-fold longer prediction horizons then is currently anticipated based on analysis of positive Lyapunov exponents for this system, even in very chaotic model regimes. It is also demonstrated that when applied for a lossy compression of the Lorenz-96 output data in various forcing regimes, EOMC achieves several orders of magnitude smaller compression loss - when compared to the common PCA-related linear compression approaches that build a backbone of the state-of-the-art lossy data compression tools (like JPEG, MP3, and others). These findings open new exciting opportunities for EOMC and transfer operator theory, by improving predictive skills and performance of data-driven tools in fluid mechanics and geosciences applications.
1 Introduction
Unsupervised dimensionality reduction and manifold learning are powerful techniques used to extract informative, low-dimensional representations from high-dimensional data, often for visualisation purposes, as a preprocessing step in machine learning pipelines, and in data compression algorithms [1, 2, 3, 4, 5, 6]. Linear methods, such as Principal Component Analysis (PCA), assume that the data lies on or near a linear subspace, and analytically-exactly minimize the sum of squared Euclidean or kernelized distances between the original full-dimensional data-points and their projections on a linear manifold [7]. In contrast, nonlinear manifold learning algorithms like t-distributed Stochastic Neighbor Embedding (t-SNE), Uniform Manifold Approximation and Projection (UMAP), and methods based on diffusion and Laplacian eigenmaps ideas assume the data resides on a complex, curved manifold embedded within the high-dimensional space, and attempt to uncover this intricate local or global topological structure [3, 8, 9, 10]. While linear methods are often easier to interpret and apply, nonlinear methods are better at revealing complex patterns and clusters in data with nonlinear relationships, but can be more challenging to parameterize and interpret.
The primary distinction in computational scaling lies in the growth rate relative to the number of data points, $T$, and the original dimensionality, DD.ăPrincipal Component Analysis (PCA) is a linear method that is highly efficient and typically scales as $\mathcal{O}(TD^{2}+D^{3})$, or sometimes as $\mathcal{O}(T^{2}D)$ depending on the implementation [7]. This makes it very fast and suitable for large datasets. PCA intrinsically produces an explicit, linear mapping function (the principal components) which can be directly applied to project new, unseen data points into the low-dimensional space without re-training.ăMethods based on the eigenmap ideas and t-distributed Stochastic Neighbour Embedding (t-SNE) have a significantly higher base polynomial complexity due to the need to compute pairwise similarities in the high-dimensional space. The calculation of the initial distance matrix involves a complexity of $\mathcal{O}(DT^{2})$ [8, 3]. The optimization step in the original algorithms scales around $\mathcal{O}(T^{2})$ or even $\mathcal{O}(T^{3})$. Optimized versions, like Barnes-Hut t-SNE, improve the optimization complexity to $\mathcal{O}(DT\log(T))$ or $\mathcal{O}(T\log(T))$ if the distance matrix is precomputed.ăUniform Manifold Approximation and Projection (UMAP) was designed to be more computationally efficient than t-SNE and generally scales as $\mathcal{O}(DT\log(T))$ or $\mathcal{O}(T\log(T))$ in practice [11].ăRegarding the explicit mapping function for unseen data points, eigenmap methods, t-SNE and UMAP, in their standard, non-parametric forms, inherently do not yield a generalizable function. In contrast to the linear methods like PCA that provide explicit algebraic rules for manifold projection, nonlinear methods typically only provide the specific embedding coordinates for the data points included in the training set.ă
Researchers have developed various methods for out-of-sample (OOS) extension of nonlinear manifold learning methods, but these often introduce new computational deficits and complexities compared to PCA’s simple linear projection:
  • ăInterpolation Methods: Projecting a single new point requires a search operation scaling as $\mathcal{O}(DT)$ naively, or $\mathcal{O}(D\log(T))$ using optimized search trees. The training is usually integrated into the original algorithm’s runtime and does not add a separate substantial cost that in the best case remains $\mathcal{O}(DT\log(T))$ as in the baseline algorithm [12].
  • Kernel Methods: These require an expensive training phase for Kernel PCA with a complexity of around $\mathcal{O}(T^{3})$ for eigenvalue decomposition. Projecting a single new data point in inference scales as $\mathcal{O}(DT)$ due to the need to compute kernel similarity against all training points [13].
  • Parametric Methods: These involve training an auxiliary, explicit function (e.g., a neural network in Parametric UMAP) to map inputs to embeddings. The training complexity of such approaches is highly variable and depends on the network architecture, but it can be substantial. Once trained, however, the inference (projection) of a new point is very efficient, typically scaling linearly with dimensionality $\mathcal{O}(D)$ for a simple forward pass.ăIt is worth noting that more complex parametric architectures, such as the transformer models, intrinsically have a polynomial scaling for their core operations [14]. For instance, the self-attention mechanism in the original transformer model scales quadratically with the sequence length (which in a data context might relate to $T$ or a sequence of tokens representing a data point). This means that while they can potentially learn very complex nonlinear mappings, their training and application complexity remains polynomial, and often more computationally demanding than simpler neural network architectures used for Parametric UMAP, particularly as $T$ or the input representation size grows [15].
Another class of nonlinear manifold learning methods is PCA-clustering, that exploits the idea of combining two linearly-scaling methods - clustering methods like K-means or Hidden Markov Models (HMMs) with PCA, exchanging the squared Euclidean distance in the K-means loss function with the squared Euclidean distance between the data points and their projections on cluster-specific linear manifolds [16, 17, 18]. As will be shown below on several examples, despite of its linear iteration complexity scaling in $T$, PCA-clustering struggles to approximate nonlinear low-dimensional manifolds.
In the following, we will start by briefly introducing the mathematical formulation of PCA-clustering (more details can be found in [16, 17, 18]), followed by demonstrating that the main bottleneck of PCA-clustering is successfully mitigated without increasing the leading order of the cost scaling, by the two modifications of its mathematical formulation: (i) by upgrading the original manifold distance loss from PCA (being a semi-norm) to the weighted combination of manifold distance and squared Euclidean distances, hereby making the clustering loss function to a weighted norm (metrisation step); (ii) and by including the Shannon entropy regularization on the cluster affiliation probability measures, making the resulting clusterings entropy-optimal (entropy regularization step). Finally, the resulting unsupervised Entropy Optimal Clustering algorithm (EOMC) will be investigated mathematically, and compared to the most popular linear and nonlinear manifold learning algorithms on synthetic benchmark examples, and on the outputs of the Lorenz-96 model in chaotic and very chaotic regimes.
2 Methods
2.1 Mathematical formulation of the PCA-clustering
Let $X\in\mathbb{R}^{T,D}$ be a $D$-dimensional real-valued data matrix with $T$ data instances, i.e., with every column $X(:,t)$ of this data matrix representing a $D$-dimensional vector of feature values for a data instance with an index $t$, where $t=1,\dots,T$ (:: denotes a column-extraction operation). Let there be $K$ clusters, each of them is characterized by its centroid $\mu_{k}\in\mathbb{R}^{D}$ and the orthogonal $d$-dimensional linear manifold projector $\mathcal{T}_{k}\in\mathbb{R}^{D,d}$, $k=1,\dots,K$. The $D$-dimensional reconstruction $X^{\textrm{rec},(k)}(:,t)$ of a given data-point $X(:,t)$ after its projection on the linear manifold $k$ (defined by $\left\{\mu_{k},\mathcal{T}_{k}\right\}$) is given as:
$X^{\textrm{rec},(k)}(:,t)$ $=$ $\mu_{k}+\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(X(:,t)-\mu_{k}\right),$ (1)
where $\dagger$ denotes a matrix transposition operation. Then, the convex reconstruction $X^{\textrm{rec}}(:,t)$ of the original data point $X(:,t)$ from its projections on all of the cluster manifolds can be computed as a convex linear combination of individual reconstructions:
$X^{\textrm{rec}}(:,t)$ $=$ $\sum_{k=1}^{K}\gamma(k,t)X^{\textrm{rec},(k)}(:,t),$ (2)
where
$\gamma(k,t)$ $\geq$ $0,\quad\textrm{and }\sum_{k=1}^{K}\gamma(k,t)=1,\quad\forall t,k.$ (3)
For a fixed $K$ the optimal values of $\gamma$ and $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ can be found as a solution of the following minimization problem:
$\left\{\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\tilde{\mathcal{L}}$
$\tilde{\mathcal{L}}$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\|X(:,t)-\sum_{k=1}^{K}\gamma(k,t)X^{\textrm{rec},(k)}(:,t)\|_{2}^{2},$
s.t. $\mathcal{T}_{k}^{\dagger}\mathcal{T}_{k}=I_{d},$ (4)
$\gamma(k,t)\geq 0,\quad\textrm{and }\sum_{k=1}^{K}\gamma(k,t)=1,\quad\forall t,k.$
Substituting (1) in (2.1) and applying the Jensen inequality, we obtain that the values of $\gamma$ and $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ can be approximated by minimizing the Jensen’s upper bound $\mathcal{L},\mathcal{L}\geq\tilde{\mathcal{L}}$ of the problem (2.1):
$\left\{\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\mathcal{L}$
$\mathcal{L}$ $=$ $\frac{1}{T}\sum_{k=1}^{K}\sum_{t=1}^{T}\gamma(k,t)\|\left(X(:,t)-\mu_{k}\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(X(:,t)-\mu_{k}\right)\|_{2}^{2},$
s.t. $\mathcal{T}_{k}^{\dagger}\mathcal{T}_{k}=I_{d},$ (5)
$\gamma(k,t)\geq 0,\quad\textrm{and }\sum_{k=1}^{K}\gamma(k,t)=1,\quad\forall t,k.$
PCA-clustering algorithm iteratively finds locally-optimal solutions of the problem (2.1) using two of its mathematical properties: (i) for a fixed $\gamma$, this problem has an analytic solution with respect to manifold parameters $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$, provided by the cluster weighted means and the dominant eigenvectors of the cluster-weighted covariance matrices; and (ii) for the fixed manifold parameters $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ and for each of the $t$, $\gamma(k^{*},t)=1$ for $k^{*}=\arg\min_{k}\|\left(X(:,t)-\mu_{k}\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(X(:,t)-\mu_{k}\right)\|_{2}^{2}$, and $\gamma(k,t)=0$ for all $k\neq k^{*}$ delivers the analytic $\gamma$-solutions of (2.1) [^1] Please note that since the optimal $\gamma$ in this solution takes only $0/1$-values, Jensen inequality becomes the Jensen equality and, hence, $\mathcal{L}=\tilde{\mathcal{L}}$. PCA-clustering starts with a random intialization of $\gamma$ and iteratively repeats these analytic steps (i) and (ii), resulting in the monotonic convergence of $\mathcal{L}$, with overall iteration complexity scaling linearly in $T$ [16, 17, 18].
However, the Lemmas 1 and 2 below gives rise to a significant mathematical difficulty, when applying PCA-clustering to approximation of low-dimensional nonlinear manifolds.
Lemma 1.
Let $\mathcal{T}$ be a real-valued $D\times d$ matrix ($D>d$) with orthonormal columns (i.e., $\mathcal{T}^{\dagger}\mathcal{T}=I_{d}$, where $I_{d}$ be the $d\times d$ identity matrix). The kernel of the operator $B=I_{D}-\mathcal{T}\mathcal{T}^{\dagger}$ is non-empty; specifically, it contains non-zero vectors.
Proof.
We analyze the operator $B=I_{D}-P$, where $P=\mathcal{T}\mathcal{T}^{\dagger}$ is the $D\times D$ orthogonal projection matrix onto the $d$-dimensional column space of $\mathcal{T}$, denoted $C(\mathcal{T})$.
To show that the kernel of $B$ is non-empty, we need to show that there exists at least one non-zero vector $v\in\mathbb{R}^{D}$ such that $Bv=0$.
$Bv=(I_{D}-P)v=0$
IvPv=0Iv-Pv=0
$Pv=v$
Hence, any vector $v$ in the kernel of $B$ must be an eigenvector of $P$ corresponding to the eigenvalue $\lambda=1$.
The set of all eigenvectors corresponding to $\lambda=1$ forms the column space of $\mathcal{T}$, $C(\mathcal{T})$. Since $P$ is an orthogonal projection onto $C(\mathcal{T})$, any vector $v\in C(\mathcal{T})$ satisfies $Pv=v$. Because the matrix $\mathcal{T}$ has $d$ orthonormal columns, the dimension of $C(\mathcal{T})$ is exactly $d$.
The kernel of $B$ is the orthogonal complement of the column space, $C(\mathcal{T})^{\perp}$. The dimension of this kernel (called the nullity of $B$) is given by the rank-nullity theorem:
$\mathrm{dim}(\mathrm{Ker}(B))=D-\mathrm{rank}(P)=D-d$
Since the problem statement assumes $D>d$, the dimension of the kernel DdD-d is a positive integer (at least 1).
A subspace with a positive dimension contains non-zero vectors. For instance, any non-zero vector that is orthogonal to every column of $\mathcal{T}$ will be in the kernel of $B$.
Therefore, the kernel of $B=I_{D}-\mathcal{T}\mathcal{T}^{\dagger}$ is non-empty and contains non-zero vectors. ∎
Lemma 2.
Let $x\in\mathbf{R}^{D}$. Then, for any $\xi\in\textrm{Ker}\left(I_{D}-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)$ and $x^{\xi}=x+\xi$ it holds that $\|\left(x-\mu_{k}\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(x-\mu_{k}\right)\|_{2}^{2}=\|\left(x^{\xi}-\mu_{k}\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(x^{\xi}-\mu_{k}\right)\|_{2}^{2}$.
Proof.
$\|\left(x^{\xi}-\mu_{k}\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(x^{\xi}-\mu_{k}\right)\|_{2}^{2}=\|\left(I_{D}-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(x^{\xi}-\mu_{k}\right)\|_{2}^{2}=$
$=\|\left(I_{D}-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(x-\mu_{k}\right)+\underbrace{\left(I_{D}-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\xi}_{\xi\in\textrm{Ker}\left(I_{D}-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)}\|_{2}^{2}=\|\left(I_{D}-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(x-\mu_{k}\right)+0\|_{2}^{2}=$
$=\|\left(I_{D}-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(x-\mu_{k}\right)\|_{2}^{2},$ (6)
where $I_{D}$ be the $D\times D$ identity matrix. ∎
From Lemmas 1 and 2 follows the first major problem of PCA-clustering: according to the Lemmas, the kernel of the projection operator is non-empty when $d<D$ (Lemma 1), and any perturbations inside of the kernel do not have any effect on the loss (Lemma 2). Hence, loss function in (2.1) is a semi-norm, not penalizing errors in the directions that are orthogonal, or close to orthogonal with respect to the manifolds $\mathcal{T}_{k}$. As will be demonstrated on examples below (see, e.g., Fig. 2C and the left panel of Fig. 3C), this can lead to the PCA-clustering minimizers $\mathcal{T}_{k}$ that become orthogonal, or close to orthogonal, with respect to the actual nonlinear low-dimensional manifold - hereby ignoring the actual manifold data, and achieving a small value of $\mathcal{L}$ at the same time. The second problem of PCA-clustering is induced by the exact $0/1$ minimizer in the step (ii) described above: it confines the resulting manifold reconstruction $X^{\textrm{rec}}(:,t)$ in (2) to piecewise-linear functions only, not allowing to achieve smooth convex interpolations of nonlinear manifolds that should be possible with the original optimization problem formulation (2.1). But, this original formulation (2.1) does not allow obtaining the cheaply-computable analytic linearly-scalable solutions (i)-(ii) that are used in the PCA-clustering algorithm. And, solving the original problem (2.1) - being highly-nonlinear, subject to multivariate and polynomial constraints - with standard numerical tools from optimization theory would require very expensive numerical algorithms that have the exponential worst-case scaling [19].
2.2 Entropy-Optimal Manifold Clustering (EOMC)
To mitigate these two central problems of PCA-clustering, the two following modifications of the loss function $\mathcal{L}$ in (2.1) will be adopted:
  • Loss metrisation: to fix the problem identified by the Lemmas 1 and 2, we will introduce an additional scalar hyper-parameter $\alpha\geq 0$ and add an addtional term $\alpha\|X(:,t)-\mu_{k}\|_{2}^{2}$ to the loss of each of the $t$ instances for all $k=1,\dots,K$. As will be proven below, this modification upgrades the semi-norm loss of (2.1) to the $\alpha$-weighted norm, fixing the problem with the data points in the directions that are close to orthogonal to $\mathcal{T}_{k}$.
  • Entropy regularization: to avoid the problem with a piecewise-linearity restriction - and to allow for the $\gamma$-solutions to be not only $0/1$ (and, to achieve this without returning back to the original formulation (2.1) that would result in exponentialy-scaling worst case numerics) - we can use the fact that the columns of $\gamma$ can be interpreted as $t$-dependent $K$-dimensional probability measures. We can add the normalized negative Shannon entropy regularization terms $\beta\gamma(k,t)\textrm{log}_{K}\left(\gamma(k,t)\right)$ (with a scalar hyper-parameter $\beta\geq 0$) to the loss, to optimally-adjust the entropies of these probability measures during learning [^2] Base $K$ of the logarithm is used to normilize the entropy values to the interval $\left[0,1\right]$. This modification is motivated by the entropic learning methods that allow training the entropy-optimal probability measures by tuning this hyper-parameter $\beta$ in the loss. The $0/1$ $\gamma$-solution of the PCA-clustering is a minimum entropy measure, and is attained when setting $\beta=0$ [20, 21, 22, 23, 24].
Making these two modifications of PCA-clustering loss and re-arranging the terms we obtain:
$\left\{\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\mathcal{L}^{\textrm{EOMC}}$
$\mathcal{L}^{\textrm{EOMC}}$ $=$ $\frac{1}{T}\sum_{k=1}^{K}\sum_{t=1}^{T}\gamma(k,t)\left[\left(X(:,t)-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}\right)+\beta\textrm{log}_{K}\left(\gamma(k,t)\right)\right],$
s.t. $\mathcal{T}_{k}^{\dagger}\mathcal{T}_{k}=I_{d},$ (7)
$\gamma(k,t)\geq 0,\quad\textrm{and }\sum_{k=1}^{K}\gamma(k,t)=1,\quad\forall t,k.$
Please note that the PCA-clustering [16, 17, 18] is a special case of (2.2), attained when setting $\alpha=0$ and $\beta=0$. Standard linear PCA is also a special case of (2.2), when selecting $K=1,\alpha=0,\beta=0$.
Before we continue, we will need to formulate and to prove two auxiliary Lemmas 3 and 4, that would in the following help us to mitigate the problem induced by the Lemmas 1 and 2 above.
Lemma 3.
Let $\mathcal{T}$ be a real-valued $D\times d$ matrix with orthonormal columns (i.e., $\mathcal{T}^{\dagger}\mathcal{T}=I_{d}$, where $I_{d}$ is the $d\times d$ identity matrix). Let $I_{D}$ be the $D\times D$ identity matrix, and let $\alpha>0$ be a positive scalar. Then, the kernel of the operator $A^{\alpha}=(1+\alpha)I_{D}-\mathcal{T}\mathcal{T}^{\dagger}$ contains only the zero vector, and operator $A^{\alpha}$ is invertible.
Proof.
We analyze the properties of the operator $A^{\alpha}=(1+\alpha)I_{D}-P$, where $P=\mathcal{T}\mathcal{T}^{\dagger}$.
The matrix $P$ is a $D\times D$ orthogonal projection matrix onto the $d$-dimensional column space of $\mathcal{T}$. A key property of any orthogonal projection matrix is that its only possible eigenvalues are 0 and $1$.
To find the kernel of $A^{\alpha}$, we seek all vectors $v\in\mathbb{R}^{D}$ such that $A^{\alpha}v=0$:
$A^{\alpha}v=((1+\alpha)I_{D}-P)v=0$
Expanding the equation gives:
$(1+\alpha)v-Pv=0$
$Pv=(1+\alpha)v$
This equation is an eigenvalue problem for the matrix $P$. For a non-zero vector $v$ to exist as a solution, the scalar $(1+\alpha)$ must be an eigenvalue of $P$.
However, we are given that $\alpha>0$. Therefore, the proposed eigenvalue $(1+\alpha)$ satisfies:
$1+\alpha>1$
Since the set of all possible eigenvalues for $P$ is restricted to $\{0,1\}$, and $1+\alpha$ is strictly greater than 1, $(1+\alpha)$ cannot be an eigenvalue of $P$. Consequently, the equation $Pv=(1+\alpha)v$ has only one solution: the zero vector, $v=0$.
Thus, the kernel of $A^{\alpha}$ contains only the zero vector, which implies that the operator $A^{\alpha}$ is invertible. ∎
Lemma 4.
Let $x$ be a column vector in $\mathbb{R}^{D}$, and let $\mathcal{T}$ be a $D\times d$ orthogonal real valued matrix, with $d$ columns forming an orthonormal basis, such that $\mathcal{T}^{\dagger}\mathcal{T}=I_{d}$. Let $I_{D}$ be the $D\times D$ identity matrix, and $\alpha>0$ be a positive scalar.
Then, the quadratic form defined by the function $f(x)=x^{\dagger}A^{\alpha}x$, where $A^{\alpha}=(1+\alpha)I_{D}-\mathcal{T}\mathcal{T}^{\dagger}$, is a strictly-convex function with a unique minimum achieved at $x=0$.
Proof.
A function $f(x)=x^{\dagger}A^{\alpha}x$ is strictly convex if and only if the matrix $A^{\alpha}$ is positive definite ($A^{\alpha}\succ 0$). A unique minimum exists if the function is strictly convex and the domain is $\mathbb{R}^{D}$, with the minimum occurring where the gradient is zero ($\nabla f(x)=2A^{\alpha}x=0$).
We analyze the matrix $A=(1+\alpha)I_{D}-\mathcal{T}\mathcal{T}^{\dagger}$. Let $P=\mathcal{T}\mathcal{T}^{\dagger}$ be the $D\times D$ orthogonal projection matrix.
To show that $A^{\alpha}$ is positive definite, we must show that $x^{\dagger}A^{\alpha}x>0$ for all non-zero vectors $x\in\mathbb{R}^{D}$.
$x^{\dagger}A^{\alpha}x=x^{\dagger}\left((1+\alpha)I_{D}-P\right)x$
$x^{\dagger}A^{\alpha}x=(1+\alpha)x^{\dagger}I_{D}x-x^{\dagger}Px$
$x^{\dagger}A^{\alpha}x=(1+\alpha)\|x\|^{2}-\|Px\|^{2}$
Here, $\|x\|^{2}$ is the squared Euclidean norm of $x$. The term $Px$ represents the orthogonal projection of $x$ onto the $d$-dimensional column space of $\mathcal{T}$, $C(\mathcal{T})$. By the properties of orthogonal projections, the norm of the projected vector is always less than or equal to the norm of the original vector: $\|Px\|^{2}\leq\|x\|^{2}$.
We can rewrite the expression as:
$x^{\dagger}A^{\alpha}x=\|x\|^{2}+\alpha\|x\|^{2}-\|Px\|^{2}$
$x^{\dagger}A^{\alpha}x=(\|x\|^{2}-\|Px\|^{2})+\alpha\|x\|^{2}$
Since $\|Px\|^{2}\leq\|x\|^{2}$, the first term $(\|x\|^{2}-\|Px\|^{2})$ is non-negative. Since $\alpha>0$, the second term $\alpha\|x\|^{2}$ is strictly positive for any $x\neq 0$.
Therefore, $x^{\dagger}A^{\alpha}x>0$ for all $x\neq 0$, which proves that the matrix $A^{\alpha}$ is positive definite.
Because $A^{\alpha}$ is positive definite, the function $f(x)=x^{\dagger}A^{\alpha}x$ is strictly convex. A strictly convex function defined on $\mathbb{R}^{D}$ has a unique global minimum. The gradient of $f(x)$ is $\nabla f(x)=2A^{\alpha}x$. Setting the gradient to zero gives $2A^{\alpha}x=0$. Since $A^{\alpha}$ is invertible (as proven in the Lemma 2), the only solution is $x=0$.
Thus, the strictly-convex function $f(x)$ has a unique minimum at $x=0$. ∎
Finally, as proven in the Theorem 1 below, optimization problem (2.2) can be solved through a sequence of analytically-solvable steps - and without exceeding the computational cost of the original PCA-clustering algorithm for (2.1), i.e., with the overall leading iteration cost scaling of $\mathcal{O}(T)$.
Theorem 1.
Let $X\in\mathbb{R}^{D,T}$, and the hyper-parameters $K\geq 1$, $d\leq K$, $\alpha>0$ and $\beta\geq 0$ are fixed. Then, the EOMC problem (2.2) has the following properties:
  1. (1.)
    Step 1 of EOMC-algorithm, analytic solutions for $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$: if $\sum_{t=1}^{T}\gamma(k,t)>0$, then for a fixed $\gamma$ that satisfies the constraints in (2.2), the solution of (2.2) is provided by:
    $\mu_{k}^{*}$ $=$ $\frac{\sum_{t=1}^{T}\gamma(k,t)X(k,t)}{\sum_{t=1}^{T}\gamma(k,t)},$ (8)
    $\mathcal{T}_{k}^{*}$ $=$ $\underset{d}{\overline{\mathrm{eigvec}}}\left(\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)\right),$ (9)
    where $\underset{d}{\overline{\mathrm{eigvec}}}\left(A\right)$ denotes an operation of computing $d$ dominant eigenvectors of A (i.e., the eigenvectors that correspond to the $d$ largest eigenvalues of the symmetric positive-semidefinite operator $A$) and putting them in the $D\times d$ orthogonal matrix column-wise. Operator $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ is defined as $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)=\frac{\sum_{t=1}^{T}\gamma(k,t)\left(X(:,t)-\mu_{k}^{*}\right)\left(X(:,t)-\mu_{k}^{*}\right)^{\dagger}}{\sum_{t=1}^{T}\gamma(k,t)}$. Cost of (8-9) computation scales as $\mathcal{O}\left(TKD\left(D+1\right)+KdD^{2}\right)$.
  2. (2.)
    Step 2 of EOMC-algorithm, analytic solutions for $\gamma(k,t)$: let $g(t)=\left(g_{1}(t),\dots,g_{K}(t)\right)$, where $g_{k}(t)=\left(X(:,t)-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}\right)$, and let $g_{\min}(t)$ be the infinum of $g(t)$ for a fixed $t$. Then, for fixed $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ and when $\beta>0$, the solution of (2.2) is given by:
    $\gamma^{*}(k,t)$ $=$ $\frac{\exp\left(-\beta^{-1}\left(g_{k}(t)-g_{\min}(t)\right)\right)}{\sum_{k=1}^{K}\exp\left(-\beta^{-1}\left(g_{k}(t)-g_{\min}(t)\right)\right)}.$ (10)
    If $\beta=0$, then the solution of (2.2) is given by $\gamma^{*}(k^{*},t)=1$ for $k^{*}=\arg\min_{k}\left(g_{k}(t)-g_{\min}(t)\right)$, and $\gamma^{*}(k,t)=0$ for all $k\neq k^{*}$. Cost of computing $\gamma^{*}$ with (8-9) scales as $\mathcal{O}\left(TK\left(D+1\right)\right)$.
  3. (3.)
    Monotonicity and convergence of EOMC-algorithm: starting with a randomly-generated $\gamma$ that satisfies constraints in (2.2), and iteratively repeating the above Step 1 and Step 2, results in monotonic decrease of the function $\mathcal{L}^{\textrm{EOMC}}$, converging to a local optimum of (2.2). The overall computational iteration cost of EOMC-algorithm scales as $\mathcal{O}\left(TK\left(D+1\right)^{2}+KdD^{2}\right)$
  4. (4.)
    Projecting a new data point on the EOMC-manifold: for $\beta>0$, nonlinear projection $Y^{\textrm{proj}}$ of any new $Y\in\mathcal{R}^{D}$ on the $d$-dimensional EOMC manifold defined by $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$, can be computed explicitly as:
    $Y^{\textrm{proj}}$ $=$ $\frac{\sum_{k=1}^{K}\exp\left(-\beta^{-1}\left(g_{k}-g_{\min}\right)\right)\left(\mu_{k}+\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(Y-\mu_{k}\right)\right),}{\sum_{k=1}^{K}\exp\left(-\beta^{-1}\left(g_{k}-g_{\min}\right)\right)}$ (11)
    where $g_{k}=\left(Y-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(Y-\mu_{k}\right)$, and $g_{\min}$ is the minimal value attained by the $K$ elements of vector $g$, and $k=1,\dots,K$. If $\beta=0$, projection is computed as $Y^{\textrm{proj}}=\mu_{k^{*}}+\mathcal{T}_{k^{*}}\mathcal{T}_{k^{*}}^{\dagger}\left(Y-\mu_{k^{*}}\right)$, where $k^{*}=\arg\min_{k}\left(g_{k}(t)-g_{\min}(t)\right)$. This computation in both situations requires at most $\mathcal{O}\left(K\left(D+1\right)\right)$ operations.
Proof.
  1. (1.)
    For fixed $X\in\mathbb{R}^{T,D}$, $\gamma$ and the hyper-parameters $K\geq 1$, $d\leq K$, $\alpha>0$ and $\beta\geq 0$, for every $\mu_{k}$, $k=1,\dots,K$, according to the Lemma1 the problem (2.2) is a strictly-convex unconstrained problem with a unique minimum
    $\sum_{t=1}^{T}\gamma(t,k)\left(X(:,t)-\mu_{k}^{*}\right)=0,$
    that after re-arranging of terms provides (8) if $\sum_{t=1}^{T}\gamma(k,t)>0$. If $\sum_{t=1}^{T}\gamma(k,t)=0$ then any value of $\mu_{k}^{*}$ is a solution. Next, we fix these $\mu_{k}^{*}$, for all $k=1,\dots,K$, and consider the solution of (2.2) with respect to $\mathcal{T}_{k}$, for fixed $X\in\mathbb{R}^{T,D}$, $\gamma$ and the hyper-parameters $K\geq 1$, $d\leq K$, $\alpha>0$ and $\beta\geq 0$. Defining the $d\times d$ matrix of Lagrange multipliers $\Lambda$, we can rewrite (2.2) in the Euler-Lagrange form
    $\tilde{\mathcal{L}}_{\Lambda}^{\textrm{EOMC}}$ $=$ $\frac{1}{T}\sum_{k=1}^{K}\sum_{t=1}^{T}\gamma(k,t)\left[\left(X(:,t)-\mu_{k}^{*}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}^{*}\right)\right]+$ (12)
    $+\sum_{i,i=1}^{d}\Lambda_{i,j}\left(\left\{I_{d}\right\}_{i,j}-\mathcal{T}^{\dagger}(:,i)\mathcal{T}(:,j)\right).$
    Taking the gradients of $\tilde{\mathcal{L}}_{\Lambda}^{\textrm{EOMC}}$ with respect to $\mathcal{T}_{k}$ and $\Lambda$, setting them to zero and re-arranging terms results in the following system of equations
    $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)\mathcal{T}_{k}$ $=$ $\Lambda\mathcal{T}_{k},$ (13)
    $\mathcal{T}_{k}^{\dagger}\mathcal{T}_{k}$ $=$ $I_{d}.$
    Substituting (13) into (12), taking trace and deploying algebraic transformations that make use of the trace operation properties, we obtain that $\Lambda$ in (13) is a diagonal matrix of dominant eigenvalues of $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ (trace of $\Lambda$ should maximize the trace of $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ after projection on the orthogonal subspace $\mathcal{T}_{k}$). Hence, $\mathcal{T}_{k}$ corresponds column-wise to the $d$ dominant eigenvectors of $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ and $\mathcal{T}_{k}^{*}=\underset{d}{\overline{\mathrm{eigvec}}}\left(\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)\right)$. Summing up the cost of computing the weighted means and weighted covariances for each cluster with the cost of computing the $d$-dominant eigenvectors of symmetric positive-semidefinite matrix $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ (that is less then a cost of full diagonalization, and allows applying iterative Krylov methods with cost scaling of $\mathcal{O}\left(KdD^{2}\right)$), we obtain the cost scaling of $\mathcal{O}\left(TKD\left(D+1\right)+KdD^{2}\right)$ for Step 1 .
  2. (2.)
    These properties follow from applying Lemma 2.1 and Lemma 2.6 in [24] to (2.2). The cost of computing Step 2 consists of $\mathcal{O}\left(TKD\right)$ (computing $K$ elements of vector $g(t)$ for all $t$), and adding the costs of $\mathcal{O}\left(TK\right)$ computations of (10) for every $t$.
  3. (3.)
    As follows from the above proofs for items (1.) and (2.), each of the Step 1 and Step 2 leads to a monotonic decrease of the function value $\mathcal{L}^{\textrm{EOMC}}$ in (2.2). It is straightforward to validate that this sequence is bounded from below with $\mathcal{L}^{\textrm{EOMC}}\geq-\beta$. Hence, since iterative repetition of these steps also generates a monotonically-decreasing sequence - and it is bounded from below with $\beta$ in $\mathcal{R}^{1}$ - this monotonically-decreasing sequence is converging to some finite $\mathcal{L}^{\textrm{EOMC}}$ if $\beta<\infty$, s.t. $\|\mathcal{L}^{\textrm{EOMC}}\|<\infty$.
  4. (4.)
    These properties and cost scaling follow directly from applying property (2.) of the Theorem to a new data point $Y\in\mathcal{R}^{D}$.
2.3 Selection of EOMC hyper-parameters $d,K,\alpha,\beta$
In comparison with the state-of-the-art nonlinear manifold learning methods like t-SNE and UMAP, EOMC relies on a much smaller set of hyper-parameters. It requires to tune only the reduced manifold dimensionality $d$, the number $K$ of locally-linear manifolds for nonlinear approximation, as well as two non-negative scalar regularization parameters $\alpha$ and $\beta$. In contrast, methods like t-SNE require tuning a much larger set of hyper-parameters, that besides of the reduced manifold dimensionality parameter $d$ and number of PCA dimensions $K$, include such model-specific adjustable parameters like perplexity, exaggeration, and Barnes-Hut tradeoff parameter. In addition, UMAP model-specific hyper-parameters include the number of neighbours, metric and its weight, and minimal distance between embedded points. Moreover, numerics of these nonlinear manifold learning methods relies on (stochastic) gradient descent algorithm - that requires tuning of multiple additional hyper-parameters that can have a very strong effect on convergence and cost of the learning phase (like learning rate schedule, batch size, cache size, etc.). As shown in the Theorem 1 in the previous section, EOMC is performed without deploying the (stochastic) gradient numerics. Instead, EOMC performs an iterative repetition of two analytic solutions in Step 1 and Step 2, not requiring additional hyper-parameter tuning.
As will be demonstrated below on practical examples (for example, see Fig. 1), tuning the EOMC hyper-parameters $d$ and $K$ can be done iteratively. To select the optimal values for $d$ and $K$, for each of the $k=1,\dots,K$ manifolds one inspects the decay of eignevalues $\Lambda^{(k)}_{1,1},\Lambda^{(k)}_{2,2},\dots,\Lambda^{(k)}_{D,D}$ for the local weighted covariances $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ in the local manifolds: as follows from the Theorem 1 above, $\sum_{i=d+1}^{D}\Lambda^{(k)}_{i,i}$ quantifies the amount of squared 2-norm loss from projecting the data, that is "assigned" to this manifold $k$, through the EOMC internal coordinates $\gamma(k,:)$. Hence, it quantifies the local quality of lossy compression, that is achieved when projecting the high-dimensional data on the local low-dimensional manifolds. Since (2.2) is an unsupervised learning problem, we can not apply most of the standard hyper-paremeter selection routines from ML and AI, like cross-validation and Bayesian hyper-paremeter tuning [25]. Instead, one can use here the hyper-parameter selection from the unsupervised regularization methods in statistics and computational science, like the $L$-curve method [26]. Alternatively, in the following examples we will adopt a nonparametric information-theoretic perspective to model selection - aiming to find the hyper-parameter combinations that lead to the models combining simplicity (measured as a high lossy compression rate, computed as a ratio between the raw data complexity and EOMC model descriptor length), and quality (measured as the relative loss of compression). In contrast to common parametric measures from information theory (like Akaike and Bayesian Information Criteria) [27] that rely on validity of parametric assumptions like Gaussianity, this non-parametric model selection procedure based on lossy compression allows a robust identification of good hyper-parameter combinations across all of the benchmarks considered below. As will also be shown on application examples below, EOMC does not require a careful and precise adjustment, since the results remain robust in the broad ranges of hyper-parameters.
3 Application examples
First, we will demonstrate several applications of the EOMC algorithm introduced above, in comparison with the state-of-the-art methods, for the synthetic noisy data examples with known low-dimensional manifold structures, and for the data produced by the Lorenz-96 model from fluid mechanics.
3.1 Synthetic examples
3.2 Example 1: nonstationary mixture of data switching between 3D ball surface and 3D torus surface in 10 dimensions with noise
Figure 1: Graphic illustration of the five phases of the EOMC data analysis pipeline. Text description is provided in the Sec. 3.2.
Figure 2: Analysis results for the nonstationary data from Example 1 (switching between two-dimensional ball and torus surface manifolds in ten dimensions with noise).
First, we will consider a non-stationary process (shown on the upper left of Fig. 1), switching between the two non-linear manifolds of very different topology: a 3D torus surface and a 3D ball surface. To make the problem more challenging for manifold learning, generated 3D data matrix is further randomly rotated in 10 dimensions and subject to a 10-dimensional Gaussian noise.
We will use this first example also to illustrate the EOMC data analysis pipeline (graphically illustrated in Fig. 1), that will be also used in the following examples:
  • Phase 1: select the hyper-parameters $d,K,\alpha,\beta$ as described in Sec. 2.3 and perform the EOMC analysis according to the Theorem 1 above. Note that the application of the standard linear PCA (left panel in Fig. 1) does not reveal any spectral gap in the eigenvalues of the covariance, and does not indicate a presence of any low-dimensional manifold. In contrast, applying EOMC with arbitrarily selected $K=15$ indicates a clear spectral gap after the second eigenvalues of the localized EOMC covariance matrices $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$. This clearly indicates that the data contains non-linear manifolds with local dimensionality $d=2$.
  • Phase 2: take the matrix of internal coordinates $\gamma$ produced in the Phase 1, and either use it directly, to reconstruct the low-dimensional representation with (11) (i.e., going straight to Phase 4), or subject $\gamma$ to further clustering, for example with the FEM-clustering method [18], that deploys the cross-entropy as a distance metric for clustering the probability measures. As can be seen from the Phase 2 illustration in the Fig. 1, this immediately uncovers the original switching process that was used in the data generation. Alternatively, as explained in the Sec. 4 below, one can omit the FEM-clustering and directly visualise the low-dimensional representation of the $\gamma$ with the standard tools of nonlinear t-SNE visualisation.
  • Phase 3: if FEM-clustering was deployed in Phase 2, in this step one brings together the manifold-projected data instances belonging to the same FEM-clusters, where projection is performed with the formula (11).
  • Phase 4: visualise and inspect the obtained manifold reconstructions. When necessary, update the hyper-parameters and return to the Phase 1.
Despite of the hyper-parameter adjustment, PCA and t-SNE fail to recover the two low-dimensional manifolds from these nonstationary data (see Fig. 2A and 2B). Fig. 2C illustrates further effects of hyper-parameter selection described in Sec. 2.3: setting both $\alpha=0$ and $\beta=0$ makes (2.2) to its special case, i.e., to PCA-clustering [16, 17, 18]. It makes visible the central problem of PCA-clustering, discussed above and induced by the Lemmas 1 and 2: the kernel of the manifold projection operator is non-empty when $d<D$ (Lemma 1), and the errors orthogonal to the manifold are not visible to the method. Therefore, the approximation in this case is rather cutting through the manifold then approximating it. Setting $\alpha$ and $\beta$ to some non-zero values mitigates this problem (see Fig. 2D). Then, increasing the number $K$ of local manifolds from 15 (which was the initial guess) to 50 results in almost perfect reconstruction of both nonlinear manifolds (see Fig. 2D). Please note that obtaining this result did not require a tedious hyper-parameter tuning, results shown in Fig. 2C were obtained from only two repetitions of the EOMC pipeline described in the Fig. 1.
3.3 Example 2: nonstationary mixture of data switching between 1D peace sign contour (planar) and 1D prism contour (in 3D), embedded and rotated in 100 dimensions with noise
Figure 3: Analysis results for the nonstationary data from Example 2 (switching between one-dimensional peace sign and prism contours manifolds in hundred dimensions with noise).
Next, we will consider a non-stationary process switching between the two non-linear one-dimensional manifolds: a 1D contour of a 3D prism and a 1D contour of the 2D peace sign, containing both piece-wise linear and nonlinear parts. To make the problem even more challenging for manifold learning (then in the previous example from Sec. 3.2), generated 3D data matrix is further randomly rotated in 100 dimensions and subject to a 100-dimensional Gaussian noise.
As in the previous example, despite of the hyper-parameter tuning, both linear PCA and t-SNE fail to find these low dimensional manifolds hidden in 100 dimensions of noisy data (see Fig. 3A and 3B). Instead, t-SNE finds a lot of clusters (see Fig. 3B), that are not present in the data, representing a clear artefact of t-SNE. Fig. 3C illustrates again the problem of PCA-clustering, and Figs. 3D and 3E show the effect of the entropic regularization term in (2.2): setting $\alpha>0$ and $\beta=0$ results in "edgy" piece-wise linear approximation of the nonlinear manifold fragments (Fig. 3D), whereas setting $\alpha>0$ and $\beta>0$ allows obtaining much better and more smooth interpolations. Like in the example 1 above, obtaining the result from the Fig. 3E did not require a tedious hyper-parameter adjustment, requiring to go through the EOMC pipeline described in the Fig. 1 only once.
3.4 Analysis of data from Lorenz-63 in chaotic, strongly-chaotic and very strongly-chaotic regimes
Model description
The Lorenz-96 (L96) model was introduced by Edward Lorenz in a 1996 paper (published later in 2005) as a simplified, yet sophisticated, "toy model" of the Earth’s atmosphere for studying the fundamental issues of predictability and chaotic dynamics in spatially extended systems [28, 29]. It mimics aspects of the mid-latitude atmosphere’s non-linear dynamics, such as advection, dissipation, and external forcing, within a computationally cheap, periodic one-dimensional domain (a latitude circle). The L96 model is widely used today as a benchmark problem for data assimilation techniques, ensemble forecasting methods, and studies on the general nature of spatiotemporal chaos [30, 31, 32, 33].
The L96 Type 1 model consists of a system of $N$ coupled ordinary differential equations (ODEs), describing the time evolution of a single scalar atmospheric quantity $X_{j}$ at $N$ equally spaced grid points around a latitude circle:
$\frac{dX_{j}}{dt}=(X_{j+1}-X_{j-2})X_{j-1}-X_{j}+F\quad\text{for }j=1,\dots,N$ (14)
Periodic boundary conditions are assumed, such that indices are taken modulo $N$ (i.e., $X_{j+N}=X_{j}$ and $X_{j-N}=X_{j}$). Variables and terms in (14) have the following meaning:
  • $X_{j}$: The value of the atmospheric quantity (e.g., temperature, vorticity) at the $j$-th grid point.
  • NN: The total number of grid points in the system (system size). Common values in literature are $N=40$.
  • $t$: Time.
  • $F$: A positive, constant external forcing parameter that drives the system.
  • $(X_{j+1}-X_{j-2})X_{j-1}$: The non-linear advection term, which conserves energy in the absence of forcing and damping.
  • $X_{j}$: A linear damping (dissipation) term.
Behaviour of the L96 model changes significantly with the forcing parameter $F$. For small values of $F$ (e.g., $F<1$), the system exhibits periodic or steady-state dynamics. As $F$ increases, the system undergoes bifurcations and transitions into chaotic regimes. A commonly studied value is $F=8$, which produces robust chaotic behaviour used frequently as a standard benchmark in predictability studies. For regimes where $F\geq 7$ (which includes $F=9$ and $F=12$ investigated below), the system is considered to be in a strong or fully turbulent chaotic state [29].
Application of EOMC to L96 output data
Figure 4: EOMC analysis results for the data from Lorenz-96 model with the external forcings $F=7,10,12$ (in rows) for the dependence between compression factor and loss as functions of reduced manifold dimensionality $d$ (first column) and number $K$ of local manifolds (second column), as well as the identified trajectories of EOMC internal coordinates $\gamma$ as functions of time (third column).
Figure 5: Panel A: heat maps of the transfer operator matrices inferred from the metastable EOMC $\gamma$ time series (partially shown in the third column of Fig. 4). Panel B: mean Markovian exit times and mean relaxation times for the Markovian transfer operators inferred for $\gamma$ that resulted from the ’EOMC analysis in a range of forcings $F$ covering the chaotic and strongly-chaotic regimes of the L96-model.
In the following, we will use the common literature setting for $N=40$, and generate long time series $X\in\mathcal{R}^{40\times 30000}$ of L96 for the three forcing regimes $F=7,9,12$, with $T=30000$ and time step $\tau=0.02$, covering the total period of 600 intrinsic time units. After rescaling according to [28, 30], this corresponds to around 3’000 Earth atmospheric days. In each of the forcing regimes we use EOMC in a broad range of hyper-parameter settings, applying the nonparametric lossy compression to find the optimal hyper-parameter values, as described above in the Sec. 2.3. As can be seen from the first and the second columns of Fig. 4, for a broad range of compression ranges, the minimal compression loss is achieved for $d=3$ in all of the forcing regimes, and for $K_{opt}$ going from $100$, over $125$ and to $150$ when $F$ goes from $7$, over $9$, and to $12$. Please note from the first and the second columns of Fig. 4 that at the same levels of lossy compression as PCA, EOMC allows achieving almost two orders of magnitude smaller relative compression losses.
We inspect the time series of intrinsic EONC coordinates $\gamma$ that were computed with these optimal hyper-parameter choices. $\gamma$ encode the probabilities of points to belong to different local manifolds $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ at different time series instances. As can be seen from the right column of Fig. 4, $\gamma$ exhibits very persistent and metastable dynamics in time, indicating relatively-long stays in each of the local linear manifolds (with $d=3$). Please note that this metastability of $\gamma$ can not be an artefact of the EOMC data analysis: as can be seen from (2.2), the EOMC loss function does not contain any terms that would enforce the persistence or metastability on $\gamma$. As a matter of fact, the value of $\mathcal{L}^{\textrm{EOMC}}$ in (2.2) is invariant with respect to any permutation of the columns of data matrix $X$. This means that the observed metastability of $\gamma$ can only be an imprint of the underlying L96 dynamics - that appears to be best described by a metastable process switching between low-dimensional (with $d=3$) locally-linear manifolds. Next, for each of the obtained $\gamma$, we compute the Markovian transition operators $P$ that describe the time evolution of $\gamma(:,t)$, by means of the exact law of the total probability, i.e., $\gamma(:,t+\tau)=P\gamma(:,t)$, where $P_{i,j}=\mathbb{P}\left[\gamma(j,t+\tau)=1|\gamma(i,t)=1\right]$ [34]. Elements $P_{i,j}$ of the transfer operator $P$ contain the probabilities of transitions from state $j$ to state $i$ in a single time step $\tau$. As demonstrated by the Fig. 5A, operators $P$ are characterized by high probabilities of staying in the states (large diagonal entries) and low probabilities of transitions to other local manifold states (very small off-diagonal entries). Next, for each of the three transfer operators computed for each of the three L96 forcing regimes, we compute the relaxation times $\tau^{rel}(P)=\frac{\tau}{1-|\lambda_{2}|}$, where $\tau$ is the time step, and $\lambda_{2}$ is the second largest (in absolute value) eigenvalue of $P$ [34]. Relaxation times $\tau^{rel}(P)$ in Markov processes measure the predictability horizons: they quantify the time it takes for the Markov processes $P$ to forget the initial condition, and to converge towards the invariant density measure [35, 36, 34]. As can be seen from the Fig. 5A, $\tau^{rel}(P)$ gradually reduces from $0.73$, over $0.61$ and to $0.57$ when $F$ goes from $7$, over $9$, and to $12$.
Next we investigate the behaviour of $\tau^{rel}(P)$ and $\mathbb{E}_{k}\left[\tau_{k}^{ex}(P)\right]$ (where $\tau^{ex}_{k}(P)=\frac{\tau}{1-P_{k,k}}$ is a Markovian mean exit time from state $k$, and $\mathbb{E}_{k}$ is the mathematical expectation over all $k=1,\dots,K$) on a more dense grid of L96 forcings $F$ between $F=7$ and $F=13$ (see Fig. 5B). Mean exit times $\tau^{ex}_{k}(P)$ quantify an average time that Markov process spends in a state $k$ before leaving it. In fluid mechanics and geosciences, predictability horizons are usually measured with the error doubling times $\tau_{d}=\frac{\ln(2)}{\lambda_{1}}$, computed from a positive leading Lyapunov exponent $\lambda_{1}$. For typical atmospheric parameters simulated by the L96 model (e.g., $N=40,F=8$), the error doubling times reported in the literature are short, roughly corresponding to the short-term forecast limits observed in real-world weather prediction models (around 2-2.5 days in atmospheric terms, corresponding to 0.42 intrinsic time units of L96 with $N=40,F=8$) [28, 30]. Larger forcings $F$ mean more chaotic behaviour and much shorter doubling times, e.g., for $N=40,F=10$ doubling time and prediction horizon is 0.28 intrinsic time units [37]. As can be seen from the red curve in Fig. 5B, average Markovian mean exit times $\mathbb{E}_{k}\left[\tau_{k}^{ex}(P)\right]$ closely match these doubling times from positive Lyapunov exponents in the literature - indicating the predictability horizons quantified with Lyapunov exponents as limits for the average times it takes for the system to leave the current low-dimensional manifold and to go somewhere else. Transfer operator description revealed by the EOMC analysis in Figs. 5 and 6 goes beyond this limit: besides allowing to measure the average time $\tau_{k}^{ex}(P)$ the dynamics spends in the local manifold $k$, transfer operator provides the transition probabilities $P_{j,k}$ to all of the other local manifold states $j\neq k$, where the dynamics can go to after leaving $k$. As revealed by the relaxation times curve $\tau^{rel}(P)$ (see the blue dotted curve in the Fig. 5B), this transfer operator description roughly doubles the prediction horizon for the system, as compared to the state of the art Lyapunov exponent description.
4 Visualising EOMC internal coordinates $\gamma$ with t-SNE
Figure 6: Nonlinear t-SNE for visualisation of the EOMC internal coordinates $\gamma$ from (2.2). Panel A: visualising $\gamma$ from the Example 1 in Sec. 3.2 (left), and $\gamma$ from the Example 2 in Sec. $3.$ Panel B: t-SNE visualisation of raw data from the L96 model described in Sec. 3.4 (left column), and for EOMC $\gamma$ (right column).
As shown above, state-of-the-art nonlinear methods like t-SNE and UMAP rely on tuning of a multitude of hyper-parameters, both method- and numerics-specific. However, in examples 1 and 2 from Sections 3.2 and 3.2, t-SNE failed to recover the low-dimensional manifolds from these nonstationary and nonlinear data (see Figs. 2B and 3B). Instead, in both of the cases it produced multiple clusters that were not present in the generated data.
Alternatively, we can deploy t-SNE to directly visualise in two or three dimensions the structures like the EOMC interpolation coefficients $\gamma$. For this, in t-SNE we can not use the common Euclidean metrics, but need to deploy a distance measure that is conform with probability measures, for example, the cross-entropy or the simmetrized Kullback-Leibler divergence. As shown in the Fig. 6A, without a particular hyper-parameter tuning (by just setting the perplexity parameter of t-SNE somewhere in the range of values between 400 and 1’200) allows fully recovering all of the original low-dimensional manifolds used in the data generation, although with some distortion. This result is a bit surprising, since it does not require taking the EOMC manifold projections from the formula (11) - and indicates that the topological structure of the manifolds is already contained in the EOMC probabilities $\gamma$.
Applying t-SNE to the raw data from the Lorenz-96 model in Sec. 3.4 reveals no pronounced clusters - although they appear to be very prominent when applying the EOMC (see Figs. 4 and 5). In contrast, without any particular hyper-parameter tuning, t-SNE with the symmetrized KL-distance reveals a very pronounced cluster structure in the $\gamma$ from EOMC. In accordance with the other results obtained in the Sec. 3.4, t-SNE reveals that the number of low-dimensional manifold clusters gradually grows with the growth of the L96 extern forcing $F$ - but the topology of these manifolds looks very similar, differing only in their position and orientation.
5 Discussion
Data-driven manifold learning and dimensionality reduction methods are constituting one of the central pillars of data analysis in many areas of science. For example, in neurosciences and bioinformatics, nonlinear methods like t-SNE and UMAP are very widely used to investigate and to visualise the high-dimensional data structures, as well as to detect clusters. However, examples provided above illustrate that t-SNE - one of the most popular tools, with over 28’000 citations according to Google Scholar - can create clusters that are not present in the data (see Figs. 2B and 3B), as well as to not detect clusters when they are actually present (see the left panels of Fig. 6B). Another problems of such methods include polynomial, or, in the best cases, $\mathcal{O}\left(T\log\left(T\right)\right)$ cost scaling with the size $T$ of data statistics - as well as necessity to tune multiple model-specific and numerics-specific hyper-parameters, and with no consensus on the optimal procedure needed to select them..
As demonstrated in the Sec. 2.2 and proven in the Lemmas 3, 4, and in the Theorem 1, Entropy Optimal Manifold Clustering (EOMC) allows mitigating these problems, resulting in computational cost scaling of $\mathcal{O}\left(T\right)$, with an explicit rule (11) to project the new data points on the manifold with cost scaling of $\mathcal{O}\left(D\right)$ in the leading order, providing a very robust learning of very nonlinear manifolds from noisy and nonstationary data, and not requiring complicated hyper-parameter adjustment (as shown on the synthetic examples the Sec. 3.2 and 3.3).
Before we discuss the results obtained for Lorenz-96 from fluid mechanics when applying the EOMC method proposed in this paper (see Sec. 3.4), we will briefly recapitulate the current knowledge regarding topology and predictability bounds in the chaotic and strongly-chaotic regimes of the Lorenz-96 model:
  • High-Dimensional Chaos: As forcing $F$ increases, the system exhibits extensive spatiotemporal chaos. The fractal dimension of the attractor grows, indicating a large number of active chaotic degrees of freedom, although this dimension density may saturate in the strong driving limit [38, 37].
  • Finite Predictability Limit: Like the real atmosphere it mimics, the L96 model in these regimes shows sensitive dependence on initial conditions, leading to a finite predictability horizon [29]. Errors grow exponentially over time, characterized by a positive leading Lyapunov exponent $\lambda_{1}$.
  • Error Doubling Time: Studies often quantify predictability horizons for a system in terms of the error doubling time $\tau_{d}=\frac{\ln(2)}{\lambda_{1}}$, computed from a positive leading Lyapunov exponent $\lambda_{1}$. For typical atmospheric parameters simulated by the model (e.g., $N=40,F=8$), the error doubling time is short, roughly corresponding to the short-term forecast limits observed in real-world weather prediction models (around 2-2.5 days in atmospheric terms, corresponding to 0.42 intrinsic time units of L96 with $N=40,F=8$) [28, 30]. Larger forcings $F$ mean more chaotic behaviour and much shorter doubling times, e.g., for $N=40,F=10$ doubling time and prediction horizon is 0.28 intrinsic time units [37]
  • Predictability Bounds: While the theoretical intrinsic predictability of the atmosphere might be around two weeks, the L96 model often reproduces practical predictability limits (e.g., 4-5 days of useful forecasts) depending on the resolution $N$ and the forcing $F$ used in the specific experiment. The strong $F\geq 7$ regimes are characterized by very rapid error growth, making accurate long-term forecasting impossible without perfect models and initial conditions [29].
However, the results obtained in the Sec. 3.4 tell a somewhat different story. EOMC analysis reveals that the internal coordinates $\gamma$ in all of the considered forcing regimes, exhibit very persistent and metastable dynamics in time, indicating relatively long stays in each of the local linear manifolds (with relatively low dimensionality $d=3$). As was analysed above, this metastability of $\gamma$ can not be an artefact of the EOMC data analysis: as can be seen from (2.2), the EOMC loss function does not contain any terms that would enforce the persistence or metastability on $\gamma$. As a matter of fact, the value of $\mathcal{L}^{\textrm{EOMC}}$ in (2.2) is invariant with respect to any permutation of the columns of data matrix $X$. This means that the observed metastability of $\gamma$ for Lorenz-96 in chaotic and very chaotic regimes can only be an imprint of the underlying L96 dynamics - that appears to be best described by a metastable process switching between low-dimensional (with $d=3$) locally-linear manifolds. Analysis from Sec. 3.4 revealed that the main effect of increasing forcing was in gradually-increasing the total number $K$ of these local low-dimensional manifolds - and in slowly decreasing mean relaxation and mean exit times, that were measured from the transfer operators inferred from EOMC variables $\gamma$ (see Fig. 5). We shown that average Markovian mean exit times $\mathbb{E}_{k}\left[\tau_{k}^{ex}(P)\right]$ closely match the doubling times from positive Lyapunov exponents in the literature (see Fig. 5B) - indicating the predictability horizons quantified with Lyapunov exponents as limits for the average times it takes for the system to leave the current low-dimensional manifold and to go somewhere else. And, it was shown that the transfer operator description revealed by the EOMC analysis in Figs. 5 and 6 allows going beyond this limit: besides allowing to measure the average time $\tau_{k}^{ex}(P)$ the dynamics spends in the local manifold $k$, transfer operator provides the transition probabilities $P_{j,k}$ to all of the other local manifold states $j\neq k$, where the dynamics can go to after leaving $k$. For Markov processes, relaxation times define the predictability horizon of the system, quantifying the time that the system requires to forget its initial condition [34]. As revealed by the relaxation times curve $\tau^{rel}(P)$ (see the blue dotted curve in the Fig. 5B), this transfer operator description roughly doubles the prediction horizon for the system, as compared to the state of the art Lyapunov exponent description.
These findings open very exciting possibilities for applying various very advanced tools from transfer operator research to the areas of fluid mechanics and geosciences. Potentially-useful approaches include methods like adaptive transfer operator sampling, milestoning, Markovian transition pathways theory and algorithms, and many others [39, 40, 41].
[…]
Availability of code Code can be provided upon a reasonable request.
[…]
Acknowledgement The author would like to thank Davide Bassetti and Tim Prokosch (both RPTU Kaiserslautern-Landau), Lukas Pospisil (VTU Ostrava), Michael Groom and Terry O’Kane (both CSIRO), Georg Gottwald (U Sydney) and the other participants of the Oberwolfach Mini-Workshop "Mathematics of Entropic AI in the Natural Sciences" in 2024, discussions with whom provided a lot of motivation for picking-up this work. Christoph Schütte and Rupert Klein (both FU Berlin) pioneered transfer operator theory and PCA-clustering that provided foundations for this work. Results on Lorenz-96 long-term predictability in chaotic regime due to the metastable manifold switching were build upon the hypothesis originally formulated by Andrew J. Majda (NYU, deceased in 2021) during the private discussion in 2010 at IPAM.This work was funded by the EU Horizons project $AI4LUNGS$ (Grant Agreement No. 101080756).